# TexSenseGAN: A User-Guided System for Optimizing Texture-Related Vibrotactile Feedback Using Generative Adversarial Network

### v4

@@ Abstract
Texture rendering has attracted significant attention as a means of creating realistic experiences in human-virtual object interactions. But in practical applications, many limited device conditions do not support the complete reproduction of spatial and temporal tactile stimuli. Different frequency components of designed vibrations can activate texture-related sensations owing to similar receptors. Therefore, we can utilize corresponding vibration signals to provide tactile feedback within the constraints of limited device environments. However, designing specific vibrations for numerous real-world materials is impractical. This study proposes a human-in-the-loop vibration generation model based on user preferences. To enable users to easily control the generation of vibration samples with large parameter spaces, we introduced an optimization model based on Differential Subspace Search (DSS) and Generative Adversarial Network (GAN). With DSS, users can employ a one-dimensional slider to easily modify the high-dimensional latent space to ensure that the GAN can generate desired vibrations. We trained the generative model using an open dataset of tactile vibration data and selected five types of vibrations as target samples for the generation experiment. Extensive user experiments were conducted using the generated and real samples. The results indicated that our system could generate distinguishable samples that matched the target characteristics. Moreover, we established a correlation between subjects’ ability to distinguish real samples and their ability to distinguish generated samples.
@@ I Introduction
Virtual realm built by information technologies has become an integral part of our daily life. Among various interaction technologies, haptics, which enables natural and intuitive tactile interaction with virtual objects and surfaces, allows users to engage with this information-rich virtual world and improves the user experience [1, 2]. In the research field of haptics, the rendering of texture of objects has been an important topic to augment the user experience in virtual reality (VR) or other application scenarios [3, 4, 5]. Texture stimuli are spatial and temporal correlated, and many studies employ parameters such as velocities, directions, and forces to establish auto-regressive models for texture modeling [6]. But there are certain requirements for this kind of presentation methods for the corresponding presentation devices, making it challenging to deploy these models on devices that are currently widespread, such as smartphones and game controllers.
Rich tactile information can be presented through vibration stimuli owing to related tactile receptors [7]. This means that even if limited conditions do not allow us to create “texture” stimuli, we can still use limited devices to generate vibration signals as “texture-related” cues. There are two methods for generating vibration stimuli: recording the actual vibration using a sensor or microphone and playing it back, or manually designing the vibration waveform. The latter is useful when direct measurement is unattainable, such as in video games or when enhancing specific tactile sensations for richer feedback. In this study, we focused on generating tactile sensations by tuning parameters in vibrotactile stimuli.
The difficulty in designing such a vibration manually is that there are several parameters to tune [3]. The diversity and abundance of tactile textures in the physical world makes it impractical to create specific models for each texture. Therefore, developing a system that can dynamically accommodate a wide range of vibration patterns while optimizing the associated parameters is highly valuable.
Several studies have been conducted on the reconstruction of vibration signals through the utilization of spectrograms derived from sensor data capturing physical surface interactions. These studies provide the optimized spectrograms that match the desired texture pattern well using generative models [4, 8]. In pursuit of a broader and more versatile spectrogram of vibrotactile signals, deep learning based generative models, particularly Generative Adversatial Network (GAN), have been enlisted. These methods can effectively optimize the objective function and yield satisfactory generative results.
We also recognize that human expert knowledge may be another crucial factor influencing the optimization, particularly in tasks related to subjective perception. Since human subjective perception may not always align with certain mathematical statistical metrics, it can sometimes be more appropriate for the user to guide the model’s generative direction for tasks related to subjective perception. Consequently, integrating human judgment and sensibility into the optimization process, a concept known as “human-in-the-loop”, becomes an appealing approach to yield results that resonate with human subjective perception, as posited in [8, 9, 10].
However, optimizing vibrotactile signals, which involves multiple parameters, presents a significant challenge for manual adjustment within the latent space. Techniques, such as Sequential Line Search (SLS) [9], enable users to manipulate multiple parameters through a one-dimensional (1-D) slider. Additionally, other interactive methods have been used to involve humans in the optimization process, such as offering different generation results for users to choose from, thereby influencing the optimization direction and generating novel textures beyond the known training dataset [8]. Considering these complexities, conventional Bayesian optimization techniques, such as SLS [9], may face limitations. Our goal is to provide users with a continuous exploration approach within the latent space, ensuring the optimization process is seamless and free from potential discontinuities.
In this study, we introduced TexSenseGAN, a user-friendly system integrating GAN and DSS for user-driven optimization, focusing on the subjective “texture-related sense” guided generative model using vibration stimuli, as shown in Fig. 1. A simplified 1-D slider interface was developed for easy exploration of high-dimensional tactile latent space. We emphasized the role of human perception in optimizing and evaluating tactile feedback, highlighting the significance of integrating subjective preferences into the design process. The experiment results show that subjects achieved a 58% accuracy rate in distinguishing the categories of generated vibrations in a five-class task, compared to 74% for real vibrations. We used a linear regression model to demonstrate a significant relationship between these two groups of accuracy rates, indicating our system’s capability to recreate distinct vibration characteristics based on user preferences.
We proposed a technique for optimizing tactile parameters, enabling control over high-dimensional feature spaces. By integrating various network architectures, we created a generative model tailored specifically for optimizing tactile parameters. Despite the extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We believe this work will contribute to the development of more intuitive and realistic haptic feedback systems.
@@ II-B Generative Neural Network
Autoencoders have been used in some studies for compressing and reconstructing tactile signals owing to their straightforward architecture and unsupervised training approach [16, 17, 18]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it makes operations, such as optimization, difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
Generative Adversarial Networks (GANs), a notable generative model, are widely recognized for their success in image generation and are increasingly applied to tactile reconstruction tasks [8, 4, 19]. A GAN consists of two main components: the generator and discriminator. The generator produces samples from a specified distribution, often Gaussian noise, while the discriminator’s role is to distinguish between generated samples and real ones. Through iterative training, the discriminator’s ability to differentiate between genuine and generated samples improves, which in turn enhances the generator’s capability to produce increasingly realistic samples.
@@ II-C Human-in-the-loop Optimization
While GANs excel at generating samples that mathematically resemble real data, in some human evaluation conditions they do not inherently account for subjective cognition. In particular, we need to determine whether a small loss value equates to a similar perception or cognition. Basic computational metrics can highlight statistical differences between generated and real samples but fail to capture human subjective perception of quality. Research has been conducted to improve GAN architectures and use better loss functions to track training progress, reducing the need for developers to stare at generated samples to figure out failure modes [21]. However, the study acknowledges that they do not provide a full quantitative evaluation of generative models. More broadly, the challenge of quantitatively evaluating GAN performance is still an open problem, with no standard metrics agreed upon [22]. This highlights a gap between statistical measures of ”realness” and human subjective judgment, especially in tasks like image and audio generation where human perception still plays a key role in evaluation. Consequently, various studies have focused on optimization models incorporating a human-in-the-loop mode, which integrates human involvement into the iterative optimization process. This approach utilizes human expertise to guide the optimization direction.
@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. The G5 and G7 groups were excluded from the study, as the original stimuli were found to evoke a tactile sensation that was similar to that of other groups [13]. While more data samples could enhance the generative model’s learning, highly complex data can also hinder model convergence. Additionally, since we used an ACGAN-like auxiliary classifier to control the generation process, using the full LMT108 dataset would require the classifier to learn a challenging 108-class classification. Although CGAN could incorporate conditional information without the need for an auxiliary classifier, after comparing its generative performance (discussed further in Sec. V-B), we finally chose to train the model using a subset of the dataset’s categories. We selected two representative types of textures from each group and built a training dataset containing fourteen classes, to ensure the model could learn a sufficiently diverse range of features, thus enhancing its capacity for generalized knowledge. The selected classes are shown in Table I.
@@ IV-C Initialization
Figure 4: The vibrotactile presenting device in this research. The vibrator was placed in a sponge block as shown in (a), and a piece of acrylic board was attached to the top of the vibrator for users to touch, as shown in (b).
@@ IV-D User Interface and Presenting Device
Fig. 4 and Fig. 5 show the vibrotactile display for evaluating the designed vibration and the user interface of DSS. The device [13] was composed of a vibro-transducer (Acouve Vp210), a power amplifier (BOSE 1705II), sponge and an acrylic cover board. The vibrator was set in the sponge block to ensure that the vibration can be transmitted effectively. An acrylic plate was attached to the vibrator, with tape covering the upper surface of the acrylic to increase friction. This setup allowed participants’ fingers to make closer contact with the vibrating acrylic plate, helped them perceive the vibrations.
@@ IV-E Stage I: Generation
To assess the generative performance of our TexSenseGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TexSenseGAN can generate new samples different from the training data, we selected classes differ from the training set as generation targets. Although we following previous research used 7 groups in the training process, our further observed that G1 and G9 exhibited excessively weak vibration intensities or significant similarity to other categories from the subjective feeling. Consequently, we also excluded G1 and G9. Finally, following stratified sampling principles, we randomly selected one class from each of the remaining groups for the task. The selected classes are shown in Fig. 6.
@@
Figure 6: The selected 5 target classes to generate.
After subjects familiarized themselves with the system and characteristics of vibrations, we proceeded to the generation experiment process. Subjects were required to generate a sample corresponding to each target vibration. The first step was the initialization of the optimization process. In this step, subjects needed to use the options “Bad”, “So-so”, and “Good” to select a suitable initial value for the optimizer. During this process, subjects were only required to make rough comparisons; thus, they were instructed not to focus on detailed perceptions, but to perform quick and multiple selections, with the aim of covering an exploration range through multiple attempts. It is important to note that the three options here do not represent the actual quality of the samples. For instance, “Good” does not mean that the sample is sufficiently close to the target sample, but rather that it possesses certain characteristics similar to the target (e.g., similar frequency components), making it a suitable initial point for optimization. After entering the optimization process, subjects were encouraged to make numerous selections rather than sticking to a single iteration’s slider position selection. However, in this phase, subjects were asked to carefully perceive various aspects of vibrations, such as frequency and amplitude. We introduced several effective ways to perceive signal changes and adjustment strategies to help participants better understand the generation process and complete the task more easily. Like participants can focus on adjusting one parameter at a time, for example, the frequency, and then other aspects. This makes the changes during the adjustment process more noticeable. Participants are also advised to carefully adjust the sliders to avoid generating distorted vibration signals. Distorted signals typically indicate that the exploration has exceeded the model’s acceptable latent space, which can lead to generation failure. Therefore, such outcomes should be avoided as much as possible. Moreover, to distinguish between different slider positions, this optimization process might be more time-consuming. When consecutive iterations did not produce vibrations significantly different along the slider direction, subjects could adopt several strategies:
@@
Following strategy 1 or strategy 2 for several consecutive iterations usually allowed the optimizer to move away from the current state. During the experiment process, subjects were asked to freely choose strategies based on their subjective judgment. However, we encourage participants to primarily opt for strategies 1 or 2 and minimize the use of the reset operation during the optimization process. If participants feel fatigued, prefer not to continue with the iterative adjustments of strategies 1 or 2, or if the generated samples become distorted, they may then choose strategy 3. Since carefully perceiving subtle differences in vibration can be mentally exhausting, and prolonged exposure may cause finger numbness, potentially affecting sensitivity, participants are allowed to take breaks at any point during the experiment to maintain accuracy in their perceptions. In our experimental observations, the reset operation was used no more than three times across all user trials.
@@ V-B Problems in Other Alternatives
Figure 10: The figure shows one of the target spectrogram for each class, along with the corresponding spectrograms generated by 3 users.
Figure 11: The figure shows the waveform of the vibrations corresponding to the spectrograms in Fig. 10
@@
We compared generative models, CGAN and ACGAN, with conditional information support. Within our task, unlike ACGAN, CGAN only requires conditional information as input, thus avoiding the challenges of multi-class classification and allowing it to handle larger datasets. However, our tests demonstrated that this approach, where conditional information is used alongside latent vectors to control generation, was less effective than ACGAN’s direct use of an auxiliary classifier to shape the latent space distribution. Under our training framework, CGAN tended to produce extremely sparse output results or exhibited a tendency not to converge. ACGAN performed relatively better among the control group models we experimented with; however, it faced challenges when applied in the optimization process, such as DSS. ACGAN can differentiate features between different classes and control them, but the uneven input vector format of label + latent vector is not conducive to the optimization process of DSS. Despite these challenges with ACGAN, we incorporated a similar auxiliary classification approach in TexSenseGAN, leveraging its strong capability for latent space control.
@@
V-C User Study
The Fig. 9 illustrates changes in subjective similarity to the target sample across three stages of the generation process: “Random” represents the initial random latent vector, “Initialized” denotes samples that have been initialized with certain target characteristics, and “Optimized” indicates the final optimized result. Light thin lines show changes in each individual trial (9 trials per class, with 3 participants generating samples for 3 target samples). The dark thick line represents the average subjective similarity for each class, with error bands indicated. Overall, the optimization process demonstrates effectiveness, supported by the benefits of the initial value selection method. Nevertheless, certain cases reveal instances of ineffective or even inverse optimization, with variations in effectiveness across different classes. Also, some generated spectrogram samples of Stage I are shown in Fig. 10, and the corresponding waveforms transformed from the spectrograms are shown in Fig. 11, to show the generation performance.
Figure 12: The average accuracy was calculated in each class.

### v5

@@ Abstract
Vibration rendering is essential for creating realistic tactile experiences in human-virtual object interactions, such as in video game controllers and VR devices. By dynamically adjusting vibration parameters based on user actions, these systems can convey spatial features and contribute to texture representation. However, generating arbitrary vibrations to replicate real-world material textures is challenging due to the large parameter space. This study proposes a human-in-the-loop vibration generation model based on user preferences. To enable users to easily control the generation of vibration samples with large parameter spaces, we introduced an optimization model based on Differential Subspace Search (DSS) and Generative Adversarial Network (GAN). With DSS, users can employ a one-dimensional slider to easily modify the high-dimensional latent space to ensure that the GAN can generate desired vibrations. We trained the generative model using an open dataset of tactile vibration data and selected five types of vibrations as target samples for the generation experiment. Extensive user experiments were conducted using the generated and real samples. The results indicated that our system could generate distinguishable samples that matched the target characteristics. Moreover, we established a correlation between subjects’ ability to distinguish real samples and their ability to distinguish generated samples.
@@ I Introduction
Virtual realm built by information technologies has become an integral part of our daily life. Among various interaction technologies, haptics, which enables natural and intuitive tactile interaction with virtual objects and surfaces, allows users to engage with this information-rich virtual world and improves the user experience [1, 2]. In the research field of haptics, the rendering of texture of objects has been an important topic to augment the user experience in virtual reality (VR) or other application scenarios [3, 4, 5]. Texture stimuli are characterized by spatially and temporally correlated signals. Many studies have modeled this kind of stimuli using parameters like velocity, direction, and force, often employing approaches such as auto-regression models [6]. Furthermore, research has shown that realistic tactile feedback can be achieved even using only temporal stimuli, such as vibrations [7, 8]. In this study, we propose a new framework for generating arbitrary vibration waveforms with the aim of presenting texture related feedback.
Currently, common vibration presenting methods include playing back vibrations recorded using sensors or microphones, and manually designing vibration waveforms through tools like equalizers. While manual design is valuable in scenarios where direct measurement is unavailable, such as video games or when enhancing specific tactile sensations, it is limited by the complexity of simultaneously adjusting multiple parameters [3], and the diversity and abundance of tactile textures in the physical world can also make this kind of adjustment impractical to create specific models for each texture.
To address these challenges, studies have explored reconstructing vibrations using variety methods like generative models [4, 9], optimizing objective functions and delivering effective results. However, human subjective perception often diverges from mathematical metrics or direct parameter adjustments, making it beneficial to let users guide the generative process for tasks involving subjective evaluation. This human-in-the-loop approach incorporates human judgment and preferences into optimization, enabling models to produce results that align with individual perceptions [9, 10, 11]. These mentioned points highlight the importance of developing a system capable of intuitively generating vibrations and dynamically accommodating a wide range of vibration patterns.
As mentioned above, adjustment of vibrotactile signals with multiple parameters manually is a challenging task. Techniques like Sequential Line Search (SLS) [10] allow users control multiple parameters via a 1-D slider simultaneously, making the process more practical. However, this kind of conventional Bayesian optimization approaches may struggle with complexities, especially in scenarios where we want to use deep learning models with a better generalization performance.
In this study, we introduced TexSenseGAN, a user-driven system combining Generative Adversarial Networks (GAN) and Differential Subspace Search (DSS) to optimize vibration stimuli for subjective “texture-related sense” generation (Fig. 1). A simple 1-D slider interface was developed to explore the high-dimensional tactile latent space easily. Emphasizing human perception, we integrated subjective preferences into optimizing and evaluating tactile feedback. The experimental results showed a 58% accuracy rate for distinguishing generated vibrations in a five-class task, compared to 74% for real vibrations. A linear regression analysis revealed a correlation between these accuracy rates, demonstrating the system’s ability to recreate distinct vibration characteristics guided by user input. Despite extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We believe this work will contribute to the development of more intuitive and realistic haptic feedback systems.
@@ II-B Generative Neural Network
Autoencoders have been used in some studies for compressing and reconstructing tactile signals owing to their straightforward architecture and unsupervised training approach [16, 17]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it makes operations, such as optimization, difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
Generative Adversarial Networks, a kind of notable generative model, are widely recognized for their success in image generation and are increasingly applied to tactile reconstruction tasks [9, 4, 18]. A GAN consists of two main components: the generator and discriminator. The generator produces samples from a specified distribution, often Gaussian noise, while the discriminator’s role is to distinguish between generated samples and real ones. Through iterative training, the discriminator’s ability to differentiate between genuine and generated samples improves, which in turn enhances the generator’s capability to produce increasingly realistic samples.
@@ II-C Human-in-the-loop Optimization
While GANs excel at generating samples that mathematically resemble real data, in some human evaluation conditions they do not inherently account for subjective cognition. In particular, we need to determine whether a small loss value equates to a similar perception or cognition. Basic computational metrics can highlight statistical differences between generated and real samples but fail to capture human subjective perception of quality. Research has been conducted to improve GAN architectures and use better loss functions to track training progress, reducing the need for developers to stare at generated samples to figure out failure modes [20]. However, the study acknowledges that they do not provide a full quantitative evaluation of generative models. More broadly, the challenge of quantitatively evaluating GAN performance is still an open problem, with no standard metrics agreed upon [21]. This highlights a gap between statistical measures of “realness” and human subjective judgment, especially in tasks like image and audio generation where human perception still plays a key role in evaluation. Consequently, various studies have focused on optimization models incorporating a human-in-the-loop mode, which integrates human involvement into the iterative optimization process. This approach utilizes human expertise to guide the optimization direction.
@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. According to a previous study, the original stimuli of G5 and G7 groups were found to evoke a tactile sensation that was similar to that of other groups [7], therefore these two groups were excluded from the study. While more data samples could enhance the generative model’s learning, highly complex data can also hinder model convergence. Additionally, since we used an ACGAN-like auxiliary classifier to control the generation process, using the full LMT108 dataset would require the classifier to learn a challenging 108-class classification. This type of multi-classification problem represents a challenging task in deep learning field. Given that it is not the primary focus of our research, we chose to avoid scenarios involving an excessively large number of categories. Although CGAN could incorporate conditional information without the need for an auxiliary classifier, after comparing its generative performance (discussed further in Sec. V-B), we finally chose to train the model using a subset of the dataset’s categories. We selected two representative types of textures from each group and built a training dataset containing fourteen classes, to ensure the model could learn a sufficiently diverse range of features, thus enhancing its capacity for generalized knowledge. The selected classes are shown in Table I.
@@ IV-C Initialization
Figure 4: The vibrotactile presenting device in this research. The vibrator was placed in a foam block as shown in (a), and a piece of acrylic board was attached to the top of the vibrator for users to touch, as shown in (b).
@@ IV-D User Interface and Presenting Device
Fig. 4 and Fig. 5 show the vibrotactile display used for evaluating the designed vibration and the user interface of the DSS. The device [7] consists of a vibro-transducer (Acouve Vp210), a power amplifier (BOSE 1705II), foam, and an acrylic cover board. The vibrator was set in the foam block to ensure that the vibration can be transmitted effectively. An acrylic plate was attached to the vibrator, and a layer of high-friction electrical tape was applied to the upper surface of the plate. This adjustment enhanced the friction, ensuring participants’ fingers maintained closer contact with the vibrating surface and enabling them to perceive the vibrations more effectively.
@@ IV-E Stage I: Generation
To assess the generative performance of our TexSenseGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TexSenseGAN can generate new samples different from the training data, we selected classes different from the training set as generation targets. Although we followed previous research and used 7 groups in the training process, our further observation was that G1 and G9 exhibited excessively weak vibration intensities or significant similarity to other categories from the subjective feeling. Consequently, we also excluded G1 and G9. Finally, following the principles of stratified sampling, we randomly selected one class from each of the remaining groups for the task. The selected classes of the LMT108 dataset are shown in Fig. 6.
@@
Figure 6: The figure shows selected 5 target classes in the LMT108 dataset to generate.
After subjects familiarized themselves with the system and characteristics of vibrations, we proceeded to the generation experiment process. Subjects were required to generate a sample corresponding to each target vibration. The first step was the initialization of the optimization process. In this step, subjects needed to use the options “Bad”, “So-so”, and “Good” to select a suitable initial value for the optimizer. During this process, subjects were only required to make rough comparisons; thus, they were instructed not to focus on detailed perceptions, but to perform quick and multiple selections, with the aim of covering an exploration range through multiple attempts. It is important to note that the three options here do not represent the actual quality of the samples. For instance, “Good” does not mean that the sample is sufficiently close to the target sample, but rather that it possesses certain characteristics similar to the target (e.g., similar frequency components), making it a suitable initial point for optimization. After entering the optimization process, subjects were encouraged to make numerous selections rather than sticking to a single iteration’s slider position selection.
However, in this phase, subjects were asked to carefully perceive various aspects of vibrations, such as frequency and amplitude. We introduced several effective ways of perceiving signal changes and adjustment strategies to help participants better understand the generation process and complete the task more easily. Participants are encouraged to adjust one parameter at a time, such as frequency, before proceeding to other aspects. This approach enhances the clarity of changes during the tuning process. To ensure optimal results, participants should handle the sliders carefully, as excessive adjustments may produce distorted vibration signals. Such distortions often indicate that the adjustments have pushed beyond the model’s acceptable latent space, potentially leading to generation failures. Therefore, avoiding these results is crucial. Moreover, to distinguish between different slider positions, this optimization process might be more time-consuming. When consecutive iterations did not produce vibrations significantly different along the slider direction, subjects could adopt several strategies:
@@
Following strategy 1 or strategy 2 for several consecutive iterations usually allowed the optimizer to move away from the current state. During the experiment process, subjects were asked to freely choose strategies based on their subjective judgment. However, we encourage participants to primarily opt for strategies 1 or 2 and minimize the use of the reset operation during the optimization process. If participants feel fatigued, prefer not to continue with the iterative adjustments of strategies 1 or 2, or if the generated samples become distorted, they may then choose strategy 3. Since carefully perceiving subtle differences in vibration can be mentally exhausting, and prolonged exposure may cause finger numbness, potentially affecting sensitivity, participants are allowed to take breaks at any point during the experiment to maintain accuracy in their perceptions. In our experimental observations, reset operation was rarely used; the subject who used the reset operation the most did so only three times throughout the entire task, and rest subjects almost never used it.
@@ IV-F Stage II: Evaluation
Figure 10: The figures show one target spectrogram for each class, along with the corresponding spectrograms generated by three users. The leftmost column represents the real target vibrations used as references, while the three columns on the right each represent one of the vibrations generated by individual users. Each row corresponds to a different vibration class.
Figure 11: The figures show the waveforms of the vibrations corresponding to the spectrograms in Fig. 10
Figure 12: The average accuracy was calculated in each class.
@@ V-B Problems in Other Alternatives
Figure 13: (a) The confusion matrix of the classification task of generated vibrations, (b) The confusion matrix of the classification task of real vibrations, (c) The average cosine similarity matrix between generated samples and target samples.
We compared generative models, CGAN and ACGAN, with conditional information support. Within our task, CGAN only requires conditional information as input, thus avoiding the challenges of multi-class auxiliary classification in ACGAN, allowing it to handle more complex datasets. However, our tests demonstrated that this approach of using conditional information alongside latent vectors as input to control generation, was less effective than ACGAN’s direct use of an auxiliary classifier to shape the latent space distribution. Under our training framework, CGAN tended to produce extremely sparse output results or exhibited a tendency not to converge. ACGAN performed relatively better among the control group models we experimented with; however, it faced challenges when applied in the optimization process, such as DSS. ACGAN can differentiate features between different classes and control them, but the uneven input vector format of label + latent vector is not conducive to the optimization process of DSS. Despite these challenges with ACGAN, we incorporated a similar auxiliary classification approach in TexSenseGAN, leveraging its strong capability for latent space control.
@@
V-C Subjective Similarity and Generation Results
Fig. 9 illustrates changes in subjective similarity to the target sample across three stages of the generation process: “Random” represents the initial random latent vector, “Initialized” denotes samples that have been initialized with certain target characteristics, and “Optimized” indicates the final optimized result. Light thin lines show changes in each individual trial (9 trials per class, with 3 participants generating samples for 3 target samples). The dark thick line represents the average subjective similarity for each class, with error bands indicated. Overall, the optimization process demonstrates effectiveness, supported by the benefits of the initial value selection method. Nevertheless, certain cases reveal instances of ineffective or even inverse optimization, with variations in effectiveness across different classes. Also, some generated spectrogram samples of Stage I are shown in Fig. 10, and the corresponding waveforms transformed from the spectrograms are shown in Fig. 11, to show the generation performance.